OpenStudy (junebug223):

Can someone walk me through this PLEASE!!! Express the equation -6a + 2b = 8 as an equivalent equation that is solved for b

OpenStudy (calculusxy):

You want to basically solve the equation in terms of b (meaning that you would want to isolate it). To do this: \(\bf\large -6a + 2b = 8\) \(\large \color{red}{\large -6a + 6a + 2b = 8 + 6a}\) \(\bf \large 2b = 8 + 6a\) \(\large \color{red}{\large \frac{2b}{2} = \frac{8 + 6a}{2}}\) \(\bf \large b = 4 + 3a\)

OpenStudy (calculusxy):

This is equivalent to -6a + 2b because if you multiply \(b = 4 + 3a\) by 2 you get \(2b = 8 + 6a\). Bring \(6a\) to the other side of the equation, and you should get \(-6a + 2b = 8\).